OpenStudy (anonymous):

please help MEDALS WILL BE GIVEN which one is greater than 1

OpenStudy (anonymous):

choices \[a : \frac{ 2 }{ 3 } - b = 3 \frac{ 1 }{ 3 }\] \[b : \frac{ 1 }{ 3 } = 2 \frac{ 5 }{ 6 } -b\] \[c : \frac{ 3 }{ 4 } - b = 1\frac{ 1 }{ 8 }\] \[d : \frac{ 1 }{ 4 } -b = \frac{ 1 }{ 2 }\]

OpenStudy (skullpatrol):

What have you tried?

OpenStudy (anonymous):

i have tried a i think

OpenStudy (anonymous):

but i dont understand how i can subtract b

OpenStudy (anonymous):

can you show me what to do so that i can try to do the other one??

OpenStudy (anonymous):

can you help me please @geerky42

geerky42 (geerky42):

Isolate b.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but would i subtract two-thirds??

OpenStudy (anonymous):

ok so a is 2.66

OpenStudy (anonymous):

ok so b is 2 1/2

OpenStudy (anonymous):

it is c right

OpenStudy (anonymous):

oh wait that isn't correct